Suppose Ghana produces cocoa and timber, and the following hypothetical data is provided for 2016 and 2017.

2016 2017
Good Quantity Price Quantity Price
Cocoa 1000 2 1100 3
Timber 500 10 525 8

Between 2016 and 2017, nominal GDP grew by ...
[1] 5.25%
[2] 10%
[3] 5%
[4] 7.14%
1
Expert's answer
2019-09-23T09:05:46-0400

[4] 7.14%


Nominal GDP is simply equal to the sum of the current year price×current year quantity of all the goods.

2016: (1000×2) + (500×10) = + 2,000 + 5,000 = $ 7,000.

2017: (1100×3) + (525×8) = + 3,300 + 4,200 = $ 7,500.

nominal GDP grew by:

7,500 / 7,000 × 100 - 100 = 7.14%